Calculate the Line Integral: S**.dř = f(x,y,z)|8c3 -4 0) – f(x,y,z) A(-1 0 2) A

Work of F= ji- di F. dr from A:(-1,0, 2) to B:(3,-4,0) along any curve that goes from A to B.
Answer #1

Let us consider the vector field

\small \vec{F}(x,y,z)=\left \langle xz,0,-yz \right \rangle

Now, we have to find

\small \int_C \vec F \cdot d\vec r

where C is the line joining A(-1,0,2) and B(3,-4,0)

So, the position vector of any point on C is given by

\small \vec r(t)=(1-t)\left \langle -1,0,2 \right \rangle+t\left \langle 3,-4,0 \right \rangle=\left \langle 4t-1,-4t,2-2t \right \rangle, where t varies from 0 to 1

So, we have, differentiating with respect to t

\small \vec r'(t)=\left \langle 4,-4,-2 \right \rangle

and,

\small \vec F(r(t))=\left \langle (4t-1)(2-2t),0,-(-4t)(2-2t) \right \rangle=\left \langle -8t^2+10t-2,0,8t-8t^2 \right \rangle

So, we have

\small \vec F(r(t))\cdot \vec r'(t)\\ \\=\left \langle -8t^2+10t-2,0,8t-8t^2 \right \rangle\cdot \left \langle 4,-4,-2 \right \rangle \\ \\=\left(-8t^2+10t-2\right)\cdot \:4+0\cdot \left(-4\right)+\left(8t-8t^2\right)\left(-2\right) \\ \\ =-16t^2+24t-8

Thus, the required integral becomes

\small \int\limits_{C}{{\vec F\centerdot d\,\vec r}} = \int_{{\,0}}^{{\,1}}{{\vec F\left( {\vec r\left( t \right)} \right)\centerdot \vec r'\left( t \right)\,dt}}=\int_0^1 \left (-16t^2+24t-8 \right )\, dt

\small =-\int _0^116t^2dt+\int _0^124tdt-\int _0^18dt = -16\left [{t^3\over 3} \right ]_0^1+\left [12t^2 \right ]_0^1-8\left [t \right ]_0^1

\small =-\frac{16}{3}+12-8=-\frac{4}{3}

Hence, we have

\small \int_A^B \vec F\cdot d\vec r = -{4\over 3}